# A Survey on All-in-One Image Restoration: Taxonomy, Evaluation and Future Trends

### v1

@@ III-C2 Mixture-of-Experts
The Mixture of Experts (MoEs) concept was first introduced in the groundbreaking 1991 paper “Adaptive Mixture of Local Experts[154]. Like ensemble methods, MoEs utilize a supervised learning approach within a framework of distinct networks. Each network, referred to as an expert, is trained to specialize in a particular subset of the data, focusing on specific regions of the input space. A gating network is responsible for selecting the most suitable expert for a given input, determining the weight assigned to each expert. Both the experts and the gating network are optimized simultaneously during training. Due to the integration of multiple expert models in the MoE architecture, each expert model can be built for different data distribution and construction patterns, thus significantly improving the professional ability of large models in various subdivisions, making MoE significantly better in handling complex tasks. With the expansion of the model scale, MoE also faces the problems of training instability and overfitting, how to ensure the generalization and robustness of the model, how to balance model performance and resource consumption, etc., waiting for the continuous optimization and improvement of large model developers.
@@
Researchers in AiOIR have observed that IR model parameters tend to be degradation-specific. For example, parameters related to one type of degradation are typically inactive when dealing with other types of degradation, and zeroing out these unrelated parameters has few impact on image restoration quality illustrated in LDR [74]. This observation aligns with the concept of conditional computation in Mixture of Experts (MoEs), where sparsity plays a key role. Applying MoEs to the field of AiOIR could lead to several improvements, including faster pretraining compared to dense models and quicker inference with the same number of parameters. Moreover, MoEs facilitate the synergistic handling of both low-level upstream tasks, such as removing weather noise, and high-level downstream tasks, such as object detection and segmentation. The following section will explore various approaches that combine MoEs with AiOIR as shown in Fig. 9, offering diverse strategies to leverage the MoE framework for improving recovered images.
Figure 9: Schematic illustration of the mixture-of-experts in AiOIR.
@@
Yang et al. [74] proposed Language-driven all-in-one adverse weather removal. This method generates degradation priors based on textual descriptions of weather conditions, which are then used to guide the adaptive selection of restoration experts through a MoE structure. It enables the model to handle diverse and mixed weather conditions without the need for specific weather-type labels, streamlining the image restoration process. WM-MoE [103] introduces a weather-aware routing mechanism (WEAR) to direct image tokens to specialized experts and employs multi-scale experts (MSE) to handle diverse weather conditions effectively. This approach leverages MoE to achieve state-of-the-art performance in adverse weather removal, enhancing both image restoration and downstream tasks like segmentation. By using a Feature Modulated Expert (FME) block and an Uncertainty-aware Router (UaR), the method achieves superior performance in image restoration while significantly reducing model parameters and inference time. The MoFME [155] framework outperforms previous approaches in both image restoration and downstream tasks, demonstrating its effectiveness and efficiency. In contrast, MEASNet [110] presents a novel multi-expert adaptive selection mechanism, which leverages both local and global image features to select the most appropriate expert models for different image restoration tasks. By balancing task-specific demands and promoting resource sharing among tasks, the approach demonstrates superior performance across multiple image degradation scenarios compared to existing methods, making it a valuable contribution to the field of image restoration.
@@ III-C3 Multimodal Model
Multi-modal tasks have become increasingly pivotal in the domain of computer vision, enabling the enrichment of visual understanding through the integration of diverse information sources. The fundamental objective of multi-modal tasks is to learn valuable potential feature representations from a multitude of modalities, such as textual captions and visual images, RGB images with supplementary components like depth or thermal images, and various forms of medical imaging data. Nevertheless, multimodal models in image restoration harness data from multiple sources to improve the fidelity and robustness of restored images. These models integrate complementary modalities to address limitations inherent in single-modality approaches, particularly in scenarios involving complex degradation, such as severe noise, blur, or low-light conditions. By leveraging different types of information, multimodal models are able to enhance the structural details, texture, and overall quality of the restored image.
@@
However, multimodal models also introduce challenges, including the increased computational complexity of handling diverse data streams and the requirement for well-aligned multimodal datasets. Additionally, the process of effectively fusing different types of data, which may have varying resolutions and characteristics, remains a significant technical hurdle. Here, we summarize a variety of approaches to multimodal AiOIR methods (e.g., Clarity-ChatGPT [150], AutoDIR [70], Instruct-IPT [116], InstructIR [62]). These involve the continuous guidance of image restoration via human language instructions, as well as the use of multimodal prompts for AiOIR, as previously mentioned. Clarity-ChatGPT is the first system that bridges adaptive image processing with interactive user feedback, which innovatively integrates large language and visual models. AutoDIR automatically detects and restores images with multiple unknown degradations by the Semantic-Agnostic Blind Image Quality Assessment (SA-BIQA). InstructIR [62] trains models using common image datasets and prompts generated using GPT-4, note that this generalizes to human-written instructions .
III-C4 Other Methods
In addition to highlighting three key improvements, we also review other methods for AiOIR. Some models benefit from iterative algorithms in the deep unfolding framework, while large-scale vision models (VLMs) like CLIP and BLIP show great promise in enhancing image restoration tasks by leveraging multimodal features. Integrating network design with pretrained Mask Image Modeling (MIM) also holds significant potential. These approaches harness semantic alignment and prior knowledge to enable robust and flexible image restoration, achieving impressive results within the AiOIR domain.
Deep Unfolding Methods. Zhang et al. [156] were the first to introduce the use of the deep unfolding framework, combining CNNs with model-based approaches for specific image restoration (IR) tasks. Considering the degradation model, the target clean image can be estimated by minimizing the energy function. Using the Half-Quadratic Splitting (HQS) algorithm [157], the equation can be decomposed into two separate subproblems, each addressing the data term and the prior term individually. The optimization is carried out by iteratively solving these subproblems in an alternating fashion. The data term subproblem reduces to a simple least squares optimization, while the prior term subproblem is addressed using a trainable CNN model. We can also use the Expectation Maximization Algorithm to model the image restoration network. DRM-IR [73] enhances the flexibility of All-In-One scenarios by introducing a reference-based, task-adaptive modeling paradigm. An advanced yet efficient AiOIR framework is developed, integrating two coupled subtasks: task-adaptive degradation modeling and model-based image restoration techniques intuitively.
@@ III-C4 Other Methods
Large-scale Vision Models. Recent works have demonstrated the potential of pretrained vision-language models (VLMs) to enhance downstream tasks using generic visual and text representations [177, 178, 179]. A classic VLM typically consists of a text encoder and an image encoder, learning aligned multimodal features from noisy image-text pairs through contrastive learning [177]. BLIP [179] improves this by eliminating noisy web data with synthetic captions. LVMs have shown robust feature representation and zero-shot transfer capabilities across various tasks. Models like CLIP [112] have demonstrated effective semantic alignment between vision and language, aiding numerous downstream tasks. Self-supervised models like DINO [113] and DINO-v2 [114] have proved effective across multiple tasks, eliminating the need for labeled data. VLMs have also gained significant traction in the domain of AiOIR. Perceive-IR [111] utilizes the semantic prior knowledge and structural information mined by a DINO-v2-based guidance module to enhance the restoration process. DA-CLIP [76] trains an additional controller that adapts the fixed CLIP image encoder to predict high-quality feature embeddings.
Mask Image Modeling. Mask Image Modeling (MIM) (e.g.,  [180, 148]) is a technique in computer vision that involves training models to predict masked portions of images based on their surrounding context inspired by Mask Language Modeling [181, 151]. This approach leverages the self-supervised learning paradigm, where the model learns to reconstruct missing parts of an image, improving its understanding of visual features and representations. MIM has been shown to be effective in various tasks, including image classification, object detection, and segmentation. The MAE [148] framework effectively employs MIM to predict hidden tokens, showcasing impressive performance and generalization capabilities across a range of downstream tasks. Meanwhile, SimMIM [180] introduced a generalized masked image modeling approach based on the Swin-ViT [182] architecture. In image restoration field, Painter [183] also leverages MIM pretraining. Qin et al. [184] introduces RAM into AiOIR, a framework designed to extract intrinsic image information from corrupted images through the use of MIM pre-training, along with a fine-tuning algorithm that facilitates the transition from masked images to fully restored ones. DyNet [185] are also trained in parallel branches to reconstruct the clean image from masked degraded inputs. For the AiOIR field, combining the exploration of the network design with pretrained MIM is of great potential.
TABLE II: Summary of implementation details of some popular AiOIR methods.
  Settings Num. Detail Degradation Datasets Methods
  multiple degradation [71] 16 Gaussian Noise, Real Noise, Rain, Fog, Snow, Raindrop, Motion Blur, Defocus Blur, Lowlight, Mixed Degradations, Under-display camera(POLED), Under-display camera(TOLED), Underwater, JPEG, Mosaicking, Moire BSD68, WED, Urban100, Kodak24, McMaster, DIV2K, Flickr2K, SIDD, Rain1400, LHP, RESIDE, , NH-HAZE, DenseHaze, Snow100k, RealSnow, Raindrop, RainDS, GoPro, RealBlur, DPDD, LOL MPerceiver [71]
multiple degradation [76] 10 Noise, Rain, Fog, Snow, Raindrop, Shadow removal, Blur, Low-light, Inpainting, JPEG DIV2K, Flickr2K, Rain100H, RESIDE, Snow100k, Raindrop, SRD, GoPro, LOL, CelebaHQ, CBSD68 DA-CLIP [76]
multiple degradation [121] 7 SR, Blur, Noise, JPEG, Rain, Haze, Low-Light DIV2K, Flickr2K MiOIR [121]
rain-haze-noise [63] 3 Rain, Haze, Noise-$\sigma 15$, Noise-$\sigma 25$, Noise-$\sigma 50$ BSD400, WED, Urban100, Rain100L, RESIDE-OTS, RESIDE-SOTS AirNet [63], PromptIR [65],PIP [72], TextpromptIR [94], NDR [96], InstructIR [62], AdaIR [78], U-WADN [186], DyNet [185], DaAIR [187], AnyIR [188], MEASNet [110], HAIR [189], Perceive-IR [111]
rain-haze-noise-blur-dark [66] 5 Rain, Haze, Noise-$\sigma 25$, Blur, Low-Light BSD400, WED, Urban100, Rain100L, RESIDE-OTS, RESIDE-SOTS, Gopro, LOL IDR [66], AdaIR [78], InstructIR [62], PIP [72], DaAIR [187], AnyIR [188], MEASNet [110], HAIR [189], Perceive-IR [111]
rain-noise-blur-dark [61] 4 Rain, Noise, Blur, Low-light SIDD, merged deraining, merged debluring,LOL ProRes [61]
rain-haze-snow [67] 3 Rain, Haze, Snow SPA+,REVIDE,RealSnow WGWS-Net [67], TKMANet [100], Art [135]
rain-haze-snow [64] 3 Raindrop, Rain+Fog, Snow All-weather (Outdoor-Rain,Snow100k,Raindrop) Transweather [64], WeatherDiff [69], AWRCP [68], AirNet [56]
rain-haze-snow [173] 3 Rain, Haze, Snow WeatherStream (176, 100 training images and 11, 400 testing images) LDR [74]
@@ IV Experiments
To facilitate a comprehensive and efficient comparison of various AiOIR methods, we begin by summarizing the key datasets, experimental setups, and evaluation metrics commonly used across different tasks. Following this, we conduct a detailed comparison of existing benchmarks across several representative image restoration tasks, such as low-light enhancement, dehazing, deblurring, image super-resolution, deraining, and desnowing. This structured approach ensures a thorough evaluation of the performance and capabilities of different AiOIR methods.
IV-A Datasets and Implementation Details
Datasets. A wide range of datasets are available for AiOIR, varying significantly in terms of image quantity, quality, resolution, and diversity. Some datasets offer paired input and target images, while others provide only ground-truth images. In the latter case, LR images are generally generated by manually. For example, BSD [162] is another classical dataset for image denoising and super-resolution. BSD100 is a classical image dataset having 100 test images proposed by Martin et al.. The dataset is composed of a large variety of images ranging from natural images to object-specific such as plants, people, food, etc. Notably, the real-world distortions are usually blind/unknown, where the distributions are different from the simple synthetic distortions. The degradation mainly falls into the following 4 categories and 15 types: Noise: Gaussian, Shot, Impulse Blur: Defocus, Glass, Motion, Zoom Weather: Snow, Frost, Fog, Bright Digital: Contrast, Elastic, Pixel, JPEG Tab. I summarizes the datasets used for different AiOIR tasks, including SR, image inpainting, deblurring, denoising, shadow removal, image desnowing, image draining and image dehazing. It is composed of the released year, the number of training samples and testing samples, and short description.
Implementation Details. We also summarize the implementation details and datasets of some AiOIR methods in Tab. II. We describe the configurations in the training process and testing process, including types, task numbers, detail degradations, datasets. We classify the types based on different experiment settings. We summarize the same experimental setup as in the original paper, where the widely adopted data augmentation techniques primarily consist of rotation and flipping operations.
Evaluation Metrics. The performance of AiOIR methods is typically evaluated using three aspects of metrics: Distortion metrics(e.g., PSNR, SSIM [36]) refer to the relationship between the restored image and the original image. A higher evaluation index indicates a greater similarity between the reconstructed image and the reference image. Perceptual metrics(e.g., FID [190], LPIPS [191]) evaluate how much the image appears like a natural image, independent of its similarity to any reference image. No-reference metrics(e.g., NIQE [192], BRISQUE [193]) are commonly based on estimating deviations from natural image statistics. In addition, there are several objective and subjective metrics that play a crucial role in measuring and comparing the performances of different AiOIR algorithms, including IL-NIQE [194], NIMA [195], CLIP-IQA [196], LOE [197], Consistency [198], PI [199], and MUSIQ [200]. In this section, we provide a detailed overview of the commonly-used metrics in AiOIR.
TABLE III: Performance Comparisons of AiOIR Models on Three Challenging Datasets.
  Method Venue Params. Dehazing Deraining Denoising on BSD68 dataset [26] Average Approach

### v2

@@ III-C2 Mixture-of-Experts
The mixture-of-experts (MoE) framework, first proposed in “Adaptive Mixture of Local Experts” [150], employs multiple specialized networks alongside a gating network that dynamically assigns weights to each expert based on the input. This architecture allows experts to focus on different data distributions, enhancing model specialization and performance in complex tasks. Researchers in AiOIR have observed that IR model parameters tend to be degradation-specific. For example, parameters related to one type of degradation are typically inactive when dealing with other types of degradation, and zeroing out these unrelated parameters has little impact on image restoration quality, as illustrated in LDR [99]. This observation aligns with the concept of conditional computation in MoEs, where sparsity plays a key role. Applying MoEs to the field of AiOIR could lead to several improvements, including faster pretraining compared to dense models and quicker inference with the same number of parameters.
@@
Yang et al. [99] proposed language-driven all-in-one adverse weather removal. This method generates degradation priors based on textual descriptions of weather conditions, which are then used to guide the adaptive selection of restoration experts. Similarly, WM-MoE [151] introduces a weather-aware routing mechanism to direct image tokens to specialized experts and employs multi-scale experts to handle diverse weather conditions. Both WM-MoE and MoFME [152] enhance IR tasks and downstream tasks like segmentation, while significantly reducing model parameters and inference time. In contrast, MEASNet [82] presents a novel multi-expert adaptive selection mechanism. This mechanism leverages both local and global image features to select the most appropriate expert models, effectively balancing task-specific demands and promoting resource sharing. More recently, MoCE-IR [153] introduces complexity experts with adaptive variable-size computational units that use a spring-inspired force mechanism to dynamically allocate experts. Building on similar principles, UniRestorer [154] construct a multi-granularity degradation set and a multi-granularity MoE restoration model. LoRA-IR [155] leverages a novel mixture of low-rank experts structure, combined with a CLIP-based degradation-guided router.
@@ III-C3 Multimodal Model
Multimodal learning has become increasingly pivotal in the domain of low-level vision [156], enabling the enrichment of visual understanding through the integration of diverse information sources. The fundamental objective of multimodal tasks is to learn potential feature representations from a multitude of modalities, such as textual captions and visual images, RGB images with supplementary components like depth or thermal images, and various forms of medical imaging data. Similarly, multimodal models in IR harness data from multiple sources to improve the fidelity and robustness of restored images. These models integrate complementary modalities to address limitations inherent in single-modality approaches, particularly in scenarios involving complex degradation. By leveraging different types of information, multimodal models enhance the structural details, texture, and overall quality of the restored image.
@@
However, multimodal models also introduce challenges, including the increased computational complexity of handling diverse data streams and the requirement for well-aligned multimodal datasets. Additionally, the process of effectively fusing different types of data, which may have varying resolutions and characteristics, remains a significant technical hurdle. Here, we summarize various approaches to multimodal AiOIR methods (e.g., Clarity-ChatGPT [147], AutoDIR [61], Instruct-IPT [95], InstructIR [94], etc). These involve the continuous guidance of IR via human language instructions, as well as the use of multimodal prompts for AiOIR, as previously mentioned. Clarity-ChatGPT is notable as the system that bridges adaptive image processing with interactive user feedback, innovatively integrating large language and visual models. AutoDIR automatically detects and restores images with multiple unknown degradations by the semantic-agnostic blind image quality assessment (SA-BIQA). InstructIR [94] trains models using common image datasets and prompts generated by GPT-4, noting that this generalizes to human-written instructions. Moreover, CMAWRNet [157] leverages structure and texture decomposition images, and $\mathrm{T}^{3}\text{-}\mathrm{DiffWeather}$ [97] uses DepthAnything [98] features as a constraint for general prompts. Recently, Perceive-IR [84] leverages multi-level quality-driven prompts categorized into three tiers of quality by image-text pairs to perceive severity levels.
III-C4 Agentic IR System
Figure 6: Illustration of Agentic IR System.
The rise of Agentic IR Systems marks a paradigm shift in handling complex real-world degradations, positioning them as a natural extension of AiOIR [158, 159]. Powered by multimodal large language models (MLLMs), these systems enable autonomous analysis, dynamic planning, and adaptive model selection as shown in Fig. 6. For instance, RestoreAgent [160] introduces modular restoration guided by degradation-aware planning, while CoR [161] proposes a step-by-step chain-of-restoration strategy, albeit more suited for non-blind scenarios. Notably, AgenticIR [158] mimics human workflows through perception, scheduling, and adaptive execution using LLM-VLM collaboration. MAIR [162] structures degradation handling via a three-stage framework based on real-world priors, reducing the search space through expert agents. HybridAgent [163] further enhances efficiency by combining fast and slow agents, enabling ambiguity resolution while minimizing error propagation. More recently, Q-Agent [159] first fine-tunes MLLMs, and uses chain-of-thought to decompose multi-degradation perception. Collectively, these agentic approaches integrate reasoning, modularity, and prior knowledge to boost robustness and flexibility in multi-degradation restoration.
III-C5 Other Methods
In addition to highlighting four key improvements, we also review other methods for AiOIR. Some models benefit from iterative algorithms in the deep unfolding networks (DUNs), and integrating the network design with pretrained mask image modeling (MIM) also holds significant potential. Meanwhile, hypernetworks with dynamically generated parameters are also highly suitable for AiOIR. The exploration of visual autoregressive modeling (VAR) as a novel paradigm for low-level vision tasks has recently extended to the field of AiOIR. The details are as follows.
Deep Unfolding Methods. Under the framework of optimization theory and deep learning, DUNs enhance IR by unfolding iterative functions into end-to-end trainable architectures. Zhang et al. [164] proposed a DUN framework that integrates CNNs with model-based priors for task-specific IR. Given a degradation model, the clean image is estimated by minimizing an energy function. The half-quadratic splitting (HQS) algorithm [165] decomposes the optimization into two subproblems: one for the data fidelity term and the other for the prior. These subproblems are alternately solved, with the data term handled via least squares and the prior term approximated using CNNs. Building on this principle, DRM-IR [60] enhances the flexibility of all-in-one scenarios by introducing a reference-based and task-adaptive modeling paradigm. It establishes an efficient AiOIR framework that jointly models task-adaptive degradation and model-based restoration. Recently, Up-restore [166] employs a prompt-driven unrolling scheme inspired by ADMM and grounded in MAP estimation, effectively combining data-driven learning with physical priors. VLU-Net [149] leverages an automatic gradient estimation strategy based on VLMs, integrating multi-level information to improve feature preservation.
@@ III-C5 Other Methods
Large-scale Vision Models. Recent works have demonstrated the potential of pretrained vision-language models (VLMs) to enhance downstream tasks using generic visual and text representations [86, 193, 194]. A classic VLM typically consists of a text encoder and an image encoder, learning aligned multimodal features from image-text pairs through contrastive learning [86]. BLIP [194] improves this by eliminating noisy web data with synthetic captions. CLIP [86] has demonstrated effective semantic alignment between vision and language, aiding numerous downstream tasks. In parallel, self-supervised models such as DINO [87] and DINO-v2 [88] have shown strong performance without requiring labeled data, further expanding the applicability of pretrained representations. VLMs have also gained significant traction in the domain of AiOIR [195, 84, 85, 83]. Perceive-IR [84] and DINO-IR [83] utilize the semantic prior knowledge and structural information mined by a DINO-based guidance module to enhance the restoration process. DA-CLIP [85] trains an additional controller that adapts the fixed CLIP image encoder to predict high-quality feature embeddings. AWRaCLe [196] employs visual in-context learning, incorporating the feature differences between the clean and degraded image pairs extracted by CLIP at the decoder.
Mask Image Modeling. Mask image modeling (MIM) (e.g.,  [197, 144, 198]) is a technique in CV that involves training models to predict masked portions of images based on their surrounding context inspired by mask language modeling [199, 200]. This self-supervised approach reconstructs missing image regions, enhancing its visual representation learning. MIM has demonstrated effectiveness in visual tasks such as classification, detection, and segmentation. The MAE [144] framework effectively employs MIM to predict hidden tokens, showcasing impressive performance and generalization capabilities across a range of downstream tasks. Qin et al. [201] are the pioneers in introducing MIM into AiOIR, shifting the focus from degradation‐type discrimination to intrinsic image information recovery. They proposed mask attribute conductance (MAC) to analyze each layer’s contribution to resolving the input integrity gap and rank layers in descending order. DyNet  [202] trains two variants (small and large) concurrently to reconstruct the clean image from masked degraded inputs. Notably, both variants share weights via an intra-network block-repetition scheme, reducing GPU hours by 50%. Most recently, Cat-AIR  [203] introduces a content-aware and task-aware router that dynamically applies self-attention to complex regions and convolution to simpler ones, using learned masks to distinguish hard from easy patches within the feature map.
Hypernetworks. Hypernetworks [204] are a class of neural networks designed to generate parameters for other networks. This enables the network to dynamically adapt to varying input patterns, enhancing its flexibility. HAIR [205] introduces data-conditioned hypernetworks into the AiOIR by incorporating the plug-and-play modules into Restormer [50] to dynamically generate transformer block weights. Zhu et al. [206] propose MWFormer for multi-weather IR. A hypernetwork is employed to extract content-independent, weather-aware features. It also generates the weights and biases for feature vectors, along with parameters for depthwise convolutions and attention modules. Furthermore, it produces weather-informed feature vectors for both weather-type identification and the guidance of pretrained weather-specific models.
Visual Autoregressive. Tian et al. [207] proposed visual autoregressive modeling (VAR), a new generation paradigm that redefines the autoregressive learning and achieves efficiency and generative performance superior to stable diffusion [208]. For AiOIR, Varformer [209] explores the multi-scale representations learned by VAR and reveals its endogenous distribution alignment priors, which transition from global color information to fine-grained details. Unlike Varformer, which utilizes intermediate VAR features to guide a separate non‐generative network, RestorerVAR [210] is generative. It fully exploits the strong priors of the pretrained VAR model by introducing degraded‐image conditioning through cross‐attention at each transformer block.
@@ IV Experiments
To facilitate a comprehensive and effective comparison of various AiOIR methods, we begin by summarizing the key datasets, experimental settings, and evaluation metrics commonly used across different tasks. Following this, we conduct a detailed comparison of existing benchmarks across several representative image restoration tasks, such as low-light enhancement, dehazing, deblurring, denoising, deraining, and desnowing. This structured schema ensures a thorough evaluation of the performance and capabilities of different all-in-one image restoration methods.
IV-A Datasets and Evaluation Metrics
Datasets. A wide range of datasets are available for AiOIR, varying significantly in terms of image quantity, quality, resolution, and diversity. Tab. I summarizes the datasets used for different AiOIR tasks. Recently, FoundIR [211] contributes a million-scale high-quality dataset for IR foundational model, highlighting the potential of AiOIR in real-world scenarios.
Evaluation Metrics. The performance of AiOIR methods is typically evaluated using three aspects of metrics: Distortion metrics(e.g., PSNR, SSIM [212]) refer to the relationship between the restored image and the original image. Perceptual metrics(e.g., FID [213], LPIPS [214]) evaluate how much the image appears like a natural image. No-reference metrics(e.g., NIQE [215], BRISQUE [216]) are commonly based on estimating deviations from natural image statistics. In addition, there are several metrics that play a crucial role in comparing the performances of different AiOIR methods, including IL-NIQE [32], NIMA [217], CLIP-IQA [218], LOE [219], Consistency [220], PI [215], and MUSIQ [221].
TABLE II: Performance Comparisons of AiOIR Models on Three Challenging Datasets.
  Method Venue & Year Params Dehazing Deraining Denoising on BSD68 dataset [174] Average Approach